OpenStudy (anonymous):

the perimeter of a rectangular is 162 m . the length is 9m more then twice the width find the dimensions. what is the length?

OpenStudy (saifoo.khan):

Width = x lenght = x +9

OpenStudy (saifoo.khan):

therefore, 162 = x + x + (x + 9) + (x+9) 162 = 4x + 18

OpenStudy (blacksteel):

Okay, we know the perimeter is twice the length plus twice the width, so lets express that as 162 = 2l + 2w. We also know that l = 2w + 9. Then 162 = 4w + 18 + 2w = 6w + 18. Then 6w = 144, so w = 24. Then l = 2*24 + 9 = 57.

OpenStudy (saifoo.khan):

solve this 4x = 162 - 18 x = 144/4 x = 36

OpenStudy (saifoo.khan):

therefore 36 + 9 = 45

OpenStudy (anonymous):

so how do i figure the length?

OpenStudy (saifoo.khan):

x + 9

OpenStudy (anonymous):

okay im lost so how do i simplify the answer?

OpenStudy (blacksteel):

Okay, I can rewrite it in a more step-by-step manner, if you want.

OpenStudy (blacksteel):

The perimeter is twice the length time twice the width. If the length is L and the width is W, the perimeter P is P = 2W + 2L. we also know that P = 162, so 2W + 2L = 162. We know that the length is twice the width + 9, so L = 2W + 9. Then we can substitute this answer into the perimeter equation to get 2W + 2(2W + 9) = 162. Simplifying this gives us 6W + 18 = 162, or 6W = 144. Dividing both sides by 6 gives us W = 24. Since we now know W, we can substitute this back into our second equation, L = 2W + 9. This gives us L = 2(24) + 9, or L = 57. (Saifoo's technique was correct, but the problem reads that the length is twice the width + 9, not the width + 9.)

OpenStudy (blacksteel):

(That should be twice the length PLUS twice the width)

OpenStudy (anonymous):

so 9 plus 9?

OpenStudy (blacksteel):

Remember, we're using L to represent the length. Above, we showed that L = 57, which is what we're solving for.

OpenStudy (blacksteel):

The twice the width + 9 part of the problem is what lets us set up the equation L = 2W + 9.

OpenStudy (anonymous):

i give up on this my brain has been doing math for 5 hours

OpenStudy (anonymous):

Mine has been doing math all day...

OpenStudy (blacksteel):

I'm a mathematician... I've been doing math for the last 8 years.